# D-LEAF: Localizing and Correcting Hallucinations in Multimodal LLMs via Layer–to-head Attention Diagnostics

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) achieve strong performance on tasks like image captioning and visual question answering, but remain prone to hallucinations, where generated text conflicts with the visual input. Prior work links this partly to insufficient visual attention, but existing attention-based detectors and mitigation typically apply uniform adjustments across layers and heads, obscuring where errors originate. In this paper, we first show these methods fail to accurately localize problematic layers. Then, we introduce two diagnostics: Layer Image Attention Entropy (LIAE) which flags anomalous layers, and Image Attention Focus (IAF) which scores attention heads within those layers. Analysis shows that LIAE pinpoints faulty layers and IAF reliably ranks heads that warrant correction. Guided by these signals, we propose Dynamic Layer-wise Entropy and Attention Fusion (D-LEAF), a task-agnostic, attention-guided method that dynamically localizes and corrects errors during inference with negligible overhead. Results show our D-LEAF delivers a 53% relative improvement on standard captioning benchmarks, and on VQA both accuracy and F1-score improve by approximately 4%, substantially suppressing hallucinations while preserving efficiency.
**footnotetext: Equal Contribution.footnotetext: Corresponding Author.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have gained increasing attention for their ability to process and integrate visual and textual information. This design allows them to achieve strong performance on a variety of vision-language tasks, such as image captioning, visual question answering, and text-to-image generation (Chen et al., 2023; Zhu et al., 2023; Liu et al., 2024a). However, MLLMs often produce content that contradicts the image or the instructions, which is known as hallucination. These inconsistencies often lead to reliability issues in practical applications, particularly in domains where accuracy and factual consistency are critical (He et al., 2023; Guo et al., 2024; Zhou et al., 2025).
Traditional strategies to mitigate hallucinations in vision-language models involve instruction fine-tuning or reinforcement learning on carefully curated datasets (Gunjal et al., 2024; Jiang et al., 2024). Although effective, these approaches are typically resource-intensive and difficult to scale. To overcome these challenges, recent research has shifted to inference‑time methods, mitigating hallucinations by enhancing semantic stability (Chen et al., 2025; Wang et al., 2025a; Tang et al., 2025) or applying contrastive decoding (Wang et al., 2024; Liang et al., 2025; Jiang et al., 2025a) techniques to adjust the distribution of the final output logits. Although these methods are more effective than training‑based algorithms, they still cannot sufficiently eliminate hallucinations and incur a higher inference latency relative to the baseline.
In addition, a deeper limitation is mechanistic: prior methods rarely identify where hallucinations arise in the attention stack. Several studies implicate over-reliance on the language stream (e.g., “anchor patterns” (Huang et al., 2024) or “textual inertia” (Liu et al., 2024b)) and respond with global adjustments that increase visual weighting (Sarkar et al., 2025; Jiang et al., 2025b). However, in practice, we find that these interventions frequently apply undifferentiated suppression across all selected attention modules. This can disrupt correctly functioning heads and thus limit hallucination reduction (see in Figure 2).
@@
To address these issues, we adopt a localize before correct strategy. We first introduce two complementary diagnostics that operate during the forward pass: (1) Layer Image-Attention Entropy (LIAE), which flags unreliable layers; and (2) Image-Attention Focus (IAF), which identifies the specific attention heads within those anomalous layers that require correction. Guided by LIAE and IAF, we propose D-LEAF (Dynamic Layer-wise Entropy and Attention Fusion), a lightweight and plug-and-play method that dynamically identifies unreliable attention components and applies selective, fused corrections to only the flagged heads, avoiding blanket suppression. As illustrated in Figure 1, when the attention entropy of a layer exceeds a dynamic threshold, D-LEAF pinpoints low focus heads and injects fused corrective signals, which suppress hallucinated content (e.g., removing the spurious phrase ‘dining table’) while preserving faithful details.
In experiments, we evaluate D-LEAF on three representative MLLM architectures across three standard multimodal hallucination benchmarks, comparing against six state-of-the-art correction methods. Results show D-LEAF consistently delivers the strongest suppression, reducing hallucinations by up to 53% versus the baseline and improving VQA accuracy and F1 by approximately 4%, with only 8% throughput drop relative to greedy decoding.These results demonstrate that D-LEAF strikes an optimal balance between factual reliability, descriptive detail, and high inference speed. Our contributions are summarized as follows:
@@
  • Second, we propose a novel method, called D-LEAF. D-LEAF is a lightweight plug-and-play method that suppresses hallucinations through layer-by-layer corrections during inference: LIAE flags problematic layers, and IAF selects heads to receive fused corrective signals.
@@ 2 Related Work
Numerous studies have examined the underlying causes of hallucinations to guide the development of more fine-grained architecture-level suppression methods. Reported factors include excessive prior knowledge of LLM (Liu et al., 2024b), insufficient attention to images (Jiang et al., 2025b; Sarkar et al., 2025; You et al., 2025), and excessive attention to summary words (Huang et al., 2024). Among these mitigation strategies, attention-head-based hallucination suppression methods show promise. For example, ASCD (Wang et al., 2025b) employs positive and negative steering as two complementary mechanisms to adapt the internal attention distributions of the model. AD‑HH (Yang et al., 2025) first identifies the heads prone to hallucination offline and then detects and suppresses these heads in real time during the model’s forward pass. In contrast, SPIN (Sarkar et al., 2025) and SVAR (Jiang et al., 2025b) indiscriminately mute a subset of heads in specific layers to force the model to focus more on visual input. However, because these approaches apply uniform corrections across all layers, they lack flexibility and can still fail to eliminate hallucinations in certain cases, as illustrated in Figure 2.
To address these issues, we conducted a systematic analysis of attention‐module behavior during the forward pass of the model and introduced LIAE for layer‐wise detection of problematic heads within each decoder module. By applying targeted per-layer corrections to these specific attention heads, our method more precisely suppresses hallucinations.
@@ 3 Understanding MLLM Hallucination
In this section, we present empirical analyses to investigate the internal mechanisms behind hallucination in MLLMs from the perspective of attention. Although prior attention-based hallucination mitigation methods have achieved promising results, we observe that they still fail to produce correct answers in semantically ambiguous scenarios, as illustrated in Figure 2. We therefore re-examine these approaches in detail. We hypothesize that indiscriminately suppression of attention heads across all layers reduces hallucinations but biases the model toward generating shorter outputs. We aim to answer these three research questions: (i) Are poorly performing heads uniformly distributed across layers?; (ii) How does head suppression reduce hallucinations?; (iii) What costs does suppression of poorly performing heads incur?
3.1 Model Forward Process
We first formalize the decoder-only forward pass used by single-stream MLLMs, where visual and textual tokens are concatenated. Let $X=[x_{1},...,x_{N}]\in\mathbb{R}^{N\times d}$ denote the input embedding. The language decoder here consists of $L$ transformer blocks. (Vaswani et al., 2017). Each block contains a multi-head attention (MHA) sublayer followed by a feed-forward network (FFN), both wrapped by residual connections. MHA begins by taking the combined text and image embedding vectors $X\in R^{N\times d}$ as input, projecting them into the query $(Q)$, key $(K)$, and value $(V)$ spaces, and then computing the attention output. The output of MHA is fed into the FFN module, and the final output of the current encoder block is obtained through residual flow, as shown in equation 1 and equation 2.
$x_{n}^{mid,l}=\sum_{h=1}^{H}\text{Attn}^{(l,h)}(X_{\leq n}^{l-1})W_{O}$ (1)
where $l$ denotes the index of transformer blocks, $h$ is the attention heads and midmid indicates the output of the MHA module. $W_{U}$ denotes the unembedding matrix used to project hidden states back to the vocabulary space, while $W_{O}$ denotes the output projection matrix of the MHA module.
@@ 3.2 Indiscriminate Correction Leads to Errors
Are poorly performing heads uniformly distributed across layers? Prior attention-correction strategies such as SPIN (Sarkar et al., 2025) and SVAR (Liu et al., 2024b) suppress the lowest-scoring heads at the intra‐layer level. We argue that ignoring inter-layer head performance can inadvertently suppress functionally correct heads, and thus fail to eliminate hallucinations. As illustrated in the middle of Figure 2, applying SPIN to LLaVA in a visually ambiguous scenario does not prevent the model’s hallucinated outputs. We hypothesize that this phenomenon arises because poorly performing attention heads are not uniformly distributed across all layers, but instead cluster within specific layers.
To further verify this hypothesis, we randomly sampled 500 images from the COCO2014 validation set and extracted the hallucinated tokens generated by the model. We rank every attention head across all layers by their SPIN scores, select the 50 lowest performing heads, and visualize their distribution to reveal how these underperformers cluster across layers. As illustrated in Figure 3(a), this distribution markedly deviates from uniformity, with the majority of layers falling below the mean. This suggests that indiscriminately modifying the lowest-scoring anomalous attention heads across all layers is not a principled or effective strategy.
Based on the above assumptions, in the context of Figure 2, we rank all attention heads globally by their SPIN scores, select the $k$ worst performing heads (using the same $k$ as SPIN) and suppress them. As shown on the right side of Figure 2, this global suppression successfully prevents the model from hallucination. These results motivate us to design a metric to dynamically localize anomalous attention heads during the model’s forward pass.
(a) Abnormal Layer
(b) Different Suppression Methods
Figure 3: (a) Distribution of abnormal attention heads across layers. (b) The impact of different suppression methods in LLaVA.
@@ 3.3 Muting low-focus heads reduces hallucinations at the cost of accuracy
In this part, we answer the questions ‘How does head suppression reduce hallucinations?’ and ‘What costs does suppression of poorly performing heads incur?’. Inspired by (Li et al., 2023a), we investigate whether there exist attention heads that can correctly capture image content. Following (Rohrbach et al., 2018), we adopt the CHAIR metrics, namely CHAIRS and CHAIRI, as defined in equation 8 and equation 9, and conduct experiments on LLaVA-7B. The detailed definitions of these metrics are provided in the Appendix B.1.
We first partition the attention heads by ranking them within each layer according to their cumulative attention over the image tokens. Based on this ranking, we suppress 15% of the heads per layer, yielding four experimental settings: (i) no intervention, (ii) suppressing the heads with the highest attention scores, (iii) suppressing the heads with the lowest attention scores, and (iv) randomly suppressing heads. We then evaluate each setting using four metrics: CHAIRS, CHAIRI, F1, and output length (Length), as illustrated in Figure 3(b).
Our results demonstrate that attention heads with high focus on image tokens indeed contribute to visual understanding, as suppressing them leads to a substantial increase in hallucination rates, approximately twice that of the baseline. Conversely, suppressing heads with low image attention can significantly reduce hallucination rates, but this comes at the cost of shorter outputs and a decrease in F1 scores. We attribute this drawback to the indiscriminate suppression of heads across all layers, as discussed above.
Moreover, our experiments reveal that modifying only a small subset of attention heads can substantially alter the model’s output, which is consistent with the findings in Kang et al. (2025a; b). Building on these observations, we hypothesize that increasing the visual focus of underperforming attention heads can reduce hallucinations, and that self-correction can be achieved by directly leveraging higher-scoring heads within the abnormal layers.
@@ 4 D-LEAF
We investigate in depth the internal mechanisms behind hallucinations in the last section. In this section, we introduce the Dynamic Layer-wise Entropy and Attention Fusion (D-LEAF) framework to mitigate hallucinations. Our method dynamically detects anomalous behaviors in the MHA modules of MLLMs during the forward pass and applies real-time corrections to the identified problematic components, thereby improving the reliability of the model’s outputs, as shown in Figure 5. We begin by introducing a novel metric Layer Image Attention Entropy (LIAE) for detecting anomalous behavior within each decoder module and describe how, once a module is flagged, we use a additional indicator Image Attention Focus (IAF) to pinpoint the exact attention heads that need to be corrected. We also present significance tests and correlation analyses for these metrics (see Appendix C for introductions to these tools). Finally, we detail the complete algorithmic workflow. We have verified the validity and effectiveness of these indicators in Appendix D.2 and replicated all analyses presented in this section on Shikra to validate the generalizability of our proposed metrics; details are provided in Appendix D.3.
@@ 4.1 Dynamic Layer Selection
Although several layer selection methods have been proposed, particularly in contrastive decoding, for instance, DeCo (Wang et al., 2024), which mixes the maximum-probability logits from a selected layer with those of the final layer, and MoLE (Liang et al., 2025), which adopts a hybrid selection strategy by comparing intermediate logits against the final layer, these approaches rely on a baseline, typically the output of the last layer. However, such a baseline is infeasible for attention-head-based corrections, since modifications to attention occur during the forward pass and each layer’s perturbation directly propagates to subsequent computations. This motivates the need for a new dynamic metric to detect abnormal layers.
Prior work has shown that, during inference in multimodal large language models (MLLMs), only a subset of attention heads are functionally engaged, and that elevated attention-head entropy is associated with a higher likelihood of hallucination (Kang et al., 2025b; Jiang et al., 2025b). Therefore, to capture the overall state of all attention heads in the current layer, we first introduce the Maximum Attention Matrix (MAM). For the $i$-th token of an MLLM output $t_{i}$, we define the MAM of the $l$-th layer:
@@
To validate the effectiveness of the metrics we proposed, we first randomly selected a subset of 500 images from the COCO 2014 validation set (Lin et al., 2014). We chose Mini-GPT4 for subsequent analysis. To show whether LIAE can significantly distinguish the differences between layers when the model generates real words and hallucinated words, we use greedy search in the decoding process of the above model to generate captions for the selected images, prompted by “Please help me describe the image in detail.” We use the ground truth annotation to identify the real and hallucinated words. We then calculated and plotted the distribution of LIAE when the model generated hallucinated words and real words.
To evaluate the significance of these metrics, which independent and non‑normally distributed across real and hallucinated tokens, we apply the Wilcoxon signed‑rank test (Wilcoxon, 1992). With $p<0.001$, in Figure 4 left, we confidently observe in these two models that hallucinated tokens exhibit significantly higher LIAE compared to real tokens.
Figure 4: The distributions of LIAE (left) and IAF (right) for hallucinated versus real words, together with the Wilcoxon signed‑rank test results in MiniGPT-4.
Accordingly, based on the above experiments, we use LIAE to localize layers that contain anomalous attention heads. Prior work also suggests that attention heads exhibit “text inertia” (Liu et al., 2024b). For completeness, we report in Appendix D.2 additional experiments that use Layer Image Attention Focus (LIAF) and a hybrid of the two (LIAS) as alternative criteria for abnormal-layer detection, together with ablations. We find that LIAE is more sensitive and achieves the best performance; therefore, we adopt LIAE as the sole metric for abnormal-layer localization.

### v2

# Tracing and Mitigating Hallucinations in Multimodal LLMs via Dynamic Attention Localization

@@ Abstract
Multimodal Large Language Models (MLLMs) achieve strong performance on tasks like image captioning and visual question answering, but remain prone to hallucinations, where generated text conflicts with the visual input. Prior work links this partly to insufficient visual attention, but existing attention-based detectors and mitigation typically apply uniform adjustments across layers and heads, obscuring where errors originate. In this paper, we first show these methods fail to accurately localize problematic layers. Then, we introduce two diagnostics: Layer Image Attention Entropy (LIAE) which flags anomalous layers, and Image Attention Focus (IAF) which scores attention heads within those layers. Analysis shows that LIAE pinpoints faulty layers and IAF reliably ranks heads that warrant correction. Guided by these signals, we propose Dynamic Layer-wise Entropy and Attention Fusion (D-LEAF), a task-agnostic, attention-guided method that dynamically localizes and corrects errors during inference with negligible overhead. Furthermore, by establishing a connection between D-LEAF and DPO, we provide theoretical justification for the effectiveness of D-LEAF. Results show our D-LEAF delivers a 53% relative improvement on standard captioning benchmarks, and on VQA both accuracy and F1-score improve by approximately 4%, substantially suppressing hallucinations while preserving efficiency.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have gained increasing attention for their ability to process and integrate visual and textual information. This design allows them to achieve strong performance on a variety of vision-language tasks, such as image captioning, visual question answering, and text-to-image generation [chen2023shikra, zhu2023minigpt, liu2024improved]. However, MLLMs often produce content that contradicts the image or the instructions, which is known as hallucination. These inconsistencies often lead to reliability issues in practical applications, particularly in domains where accuracy and factual consistency are critical [yao2025understanding, he2023geometric, zhang2025modalities, zhou2025flattery, guo2025benchmarking, zhang2024locate, cheng2024leveraging, cheng2024multi, yao2025your, yang2025fraud].
Traditional strategies to mitigate hallucinations in vision-language models involve instruction fine-tuning or reinforcement learning on carefully curated datasets [gunjal2024detecting, jiang2024hallucination, hu2024dissecting]. Although effective, these approaches are typically resource-intensive and difficult to scale. To overcome these challenges, recent research has shifted to inference‑time methods, mitigating hallucinations by enhancing semantic stability [chen2025mitigating, wang2025tpc, tang2025mitigating, jiang2025msrs, yu2025pixel] or applying contrastive decoding techniques [wang2024mllm, liang2025mole, jiang2025hicd] to adjust the distribution of the final output logits. Although these methods are more effective than training‑based algorithms, they still cannot sufficiently eliminate hallucinations and incur a higher inference latency relative to the baseline.
In addition, a deeper limitation is mechanistic: prior methods rarely identify where hallucinations arise in the attention stack. Several studies implicate over-reliance on the language stream (e.g., “anchor patterns” [huang2024opera] or “textual inertia” [liu2407paying]) and respond with global adjustments that increase visual weighting [sarkar2025mitigating, jiang2025devils]. However, in practice, we find that these interventions frequently apply undifferentiated suppression across all selected attention modules. This can disrupt correctly functioning heads and thus limit hallucination reduction (see in Fig. 2).
@@
To address these issues, we adopt a localize before correct strategy. We first introduce two complementary diagnostics that operate during the forward pass: (1) Layer Image-Attention Entropy (LIAE), which flags unreliable layers; and (2) Image-Attention Focus (IAF), which identifies the specific attention heads within those anomalous layers that require correction. Guided by LIAE and IAF, we propose D-LEAF (Dynamic Layer-wise Entropy and Attention Fusion), a lightweight and plug-and-play method that dynamically identifies unreliable attention components and applies selective, fused corrections to only the flagged heads, avoiding blanket suppression. As illustrated in Fig. 1, when the attention entropy of a layer exceeds a dynamic threshold, D-LEAF pinpoints low focus heads and injects fused corrective signals, which suppress hallucinated content (e.g., removing the spurious phrase ‘dining table’) while preserving faithful details.
To address these issues, we adopt a localize before correct strategy. We first introduce two complementary diagnostics that operate during the forward pass: (1) Layer Image-Attention Entropy (LIAE), which flags unreliable layers; and (2) Image-Attention Focus (IAF), which identifies the specific attention heads within those anomalous layers that require correction. Guided by LIAE and IAF, we propose D-LEAF (Dynamic Layer-wise Entropy and Attention Fusion), a lightweight and plug-and-play method that dynamically identifies unreliable attention components and applies selective, fused corrections to only the flagged heads, avoiding blanket suppression. As illustrated in Fig. 1, when the attention entropy of a layer exceeds a dynamic threshold, D-LEAF pinpoints low focus heads and injects fused corrective signals, which suppress hallucinated content (e.g., removing the spurious phrase ‘dining table’) while preserving faithful details.
These results demonstrate that D-LEAF strikes an optimal balance between factual reliability, descriptive detail, and high inference speed. Our contributions are summarized as follows:
@@
  • Second, we propose a novel method, called D-LEAF. D-LEAF is a lightweight plug-and-play method that suppresses hallucinations through layer-by-layer corrections during inference: LIAE flags problematic layers, and IAF selects heads to receive fused corrective signals. Furthermore, we show that D-LEAF and DPO both aim to guide the model toward generating preferred answers through a shared optimization objective.
@@ 2 Related Work
Numerous studies have examined the underlying causes of hallucinations to guide the development of more fine-grained architecture-level suppression methods. Reported factors include excessive prior knowledge of LLM [liu2407paying], insufficient attention to images [jiang2025devils, sarkar2025mitigating, you2025mitigating], and excessive attention to summary words [huang2024opera]. Among these mitigation strategies, attention-head-based hallucination suppression methods show promise. For example, ASCD [wang2025ascd] employs positive and negative steering as two complementary mechanisms to adapt the internal attention distributions of the model. AD‑HH [yangunderstanding] first identifies the heads prone to hallucination offline and then detects and suppresses these heads in real time during the model’s forward pass. In contrast, SPIN [sarkar2025mitigating] and SVAR [jiang2025devils] indiscriminately mute a subset of heads in specific layers to force the model to focus more on visual input. However, because these approaches apply uniform corrections across all layers, they lack flexibility and can still fail to eliminate hallucinations in certain cases, as illustrated in Fig. 2. To address these issues, we conducted a systematic analysis of attention‐module behavior during the forward pass of the model and introduced LIAE for layer‐wise detection of problematic heads within each decoder module. By applying targeted per-layer corrections to these specific attention heads, our method more precisely suppresses hallucinations.
@@ 3 Understanding MLLM Hallucination
In this section, we present empirical analyses to investigate the internal mechanisms behind hallucination in MLLMs from the perspective of attention. Although prior attention-based hallucination mitigation methods have achieved promising results, we observe that they still fail to produce correct answers in semantically ambiguous scenarios, as illustrated in Fig. 2. We therefore re-examine these approaches in detail. We hypothesize that indiscriminately suppression of attention heads across all layers reduces hallucinations but biases the model toward generating shorter outputs. We aim to answer these three research questions: (i) Are poorly performing heads uniformly distributed across layers?; (ii) How does head suppression reduce hallucinations?; (iii) What costs does suppression of poorly performing heads incur?
@@ 3.1 Indiscriminate Correction Leads to Errors
Are poorly performing heads uniformly distributed across layers? Prior attention-correction strategies such as SPIN [sarkar2025mitigating] and SVAR [liu2407paying] suppress the lowest-scoring heads at the intra‐layer level. We argue that ignoring inter-layer head performance can inadvertently suppress functionally correct heads, and thus fail to eliminate hallucinations. As illustrated in the middle of Fig. 2, applying SPIN to LLaVA in a visually ambiguous scenario does not prevent the model’s hallucinated outputs. We hypothesize that this phenomenon arises because poorly performing attention heads are not uniformly distributed across all layers, but instead cluster within specific layers.
Figure 3: Distribution of abnormal attention heads across layers.
To further verify this hypothesis, we randomly sampled 500 images from the COCO2014 validation set and extracted the hallucinated tokens generated by the model. We rank every attention head across all layers by their SPIN scores, select the 50 lowest performing heads, and visualize their distribution to reveal how these underperformers cluster across layers. As illustrated in Fig. 3, this distribution markedly deviates from uniformity, with the majority of layers falling below the mean. This suggests that indiscriminately modifying the lowest-scoring anomalous attention heads across all layers is not a principled or effective strategy.
Based on the above assumptions, in the context of Fig. 2, we rank all attention heads globally by their SPIN scores, select the $k$ worst performing heads (using the same $k$ as SPIN) and suppress them. As shown on the right side of Fig. 2, this global suppression successfully prevents the model from hallucination. These results motivate us to design a metric to dynamically localize anomalous attention heads during the model’s forward pass.
@@ 3.2 Muting low-focus heads reduces hallucinations at the cost of accuracy
Figure 4: The impact of different suppression methods in LLaVA.
In this part, we answer the questions ‘How does head suppression reduce hallucinations?’ and ‘What costs does suppression of poorly performing heads incur?’. Inspired by [li2023inference], we investigate whether there exist attention heads that can correctly capture image content. Following [rohrbach2018object], we adopt the CHAIR metrics, namely CHAIRI and CHAIRS, as defined in Eq. 7 and Eq. 8, and conduct experiments on LLaVA-7B. The detailed definitions of these metrics are provided in the Sec. B.1.
We first partition the attention heads by ranking them within each layer according to their cumulative attention over the image tokens. Based on this ranking, we suppress 15% of the heads per layer, yielding four experimental settings: (i) no intervention, (ii) suppressing the heads with the highest attention scores, (iii) suppressing the heads with the lowest attention scores, and (iv) randomly suppressing heads. We then evaluate each setting using four metrics: CHAIRS, CHAIRI, F1, and output length (Length), as illustrated in Fig. 4.
Our results demonstrate that attention heads with high focus on image tokens indeed contribute to visual understanding, as suppressing them leads to a substantial increase in hallucination rates, approximately twice that of the baseline. Conversely, suppressing heads with low image attention can significantly reduce hallucination rates, but this comes at the cost of shorter outputs and a decrease in F1 scores. We attribute this drawback to the indiscriminate suppression of heads across all layers, as discussed above. Moreover, our experiments reveal that modifying only a small subset of attention heads can substantially alter the model’s output, which is consistent with the findings in [kang2025see, kang2025your].
Building on these observations, we hypothesize that increasing the visual focus of underperforming attention heads can reduce hallucinations, and that self-correction can be achieved by directly leveraging higher-scoring heads within the abnormal layers.
@@ 4 D-LEAF
We investigate in depth the internal mechanisms behind hallucinations in the last section. In this section, we introduce the Dynamic Layer-wise Entropy and Attention Fusion (D-LEAF) framework to mitigate hallucinations. Our method dynamically detects anomalous behaviors in the MHA modules of MLLMs during the forward pass and applies real-time corrections to the identified problematic components, thereby improving the reliability of the model’s outputs, as shown in Fig. 24. We begin by introducing a novel metric Layer Image Attention Entropy (LIAE) for detecting anomalous behavior within each decoder module and describe how, once a module is flagged, we use a additional indicator Image Attention Focus (IAF) to pinpoint the exact attention heads that need to be corrected. We also present significance tests and correlation analyses for these metrics (see Appendix C for introductions to these tools). Finally, we detail the complete algorithmic workflow. We have verified the validity and effectiveness of these indicators in Sec. D.2 and replicated all analyses presented in this section on Shikra to validate the generalizability of our proposed metrics; details are provided in Sec. D.3.
@@ 4.1 Dynamic Layer Selection
Figure 5: LIAE Distribution across object token types in MiniGPT-4.
Although several layer selection methods have been proposed, particularly in contrastive decoding, for instance, DeCo [wang2024mllm], which mixes the maximum-probability logits from a selected layer with those of the final layer, and MoLE [liang2025mole], which adopts a hybrid selection strategy by comparing intermediate logits against the final layer, these approaches rely on a baseline, typically the output of the last layer. However, such a baseline is infeasible for attention-head-based corrections, since modifications to attention occur during the forward pass and each layer’s perturbation directly propagates to subsequent computations. This motivates the need for a new dynamic metric to detect abnormal layers.
Prior work has shown that, during inference in MLLMs, only a subset of attention heads are functionally engaged, and that elevated attention-head entropy is associated with a higher likelihood of hallucination [kang2025your, jiang2025devils]. Therefore, to capture the overall state of all attention heads in the current layer, we first introduce the Maximum Attention Matrix (MAM). For the $i$-th token of an MLLM output $t_{i}$, we define the MAM of the $l$-th layer:
@@
To validate the effectiveness of the metrics we proposed, we first randomly selected a subset of 500 images from the COCO 2014 validation set [lin2014microsoft]. We chose MiniGPT-4 for subsequent analysis. To show whether LIAE can significantly distinguish the differences between layers when the model generates real words and hallucinated words, we use greedy search in the decoding process of the above model to generate captions for the selected images, prompted by “Please help me describe the image in detail.” We use the ground truth annotation to identify the real and hallucinated words. We then calculated and plotted the distribution of LIAE when the model generated hallucinated words and real words.
To evaluate the significance of these metrics, which independent and non‑normally distributed across real and hallucinated tokens, we apply the Wilcoxon signed‑rank test [wilcoxon1992individual]. With $p<0.001$, in Fig. 5, we confidently observe in these two models that hallucinated tokens exhibit significantly higher LIAE compared to real tokens.
Accordingly, based on the above experiments, we use LIAE to localize layers that contain anomalous attention heads. For completeness, we report in Sec. D.2 additional experiments that use Layer Image Attention Focus (LIAF) and a hybrid of the two (LIAS) as alternative criteria for abnormal-layer detection, together with ablations. We find that LIAE is more sensitive and achieves the best performance; therefore, we adopt LIAE as the sole metric for abnormal-layer localization.